OpenStudy (anonymous):

Write the equation of a line with a slope of 2 and containing the point (1,4) in y=mx+b format. Please Help!

OpenStudy (anonymous):

the slope is m can you write the equation now, then i will help u with the next step

OpenStudy (anonymous):

So, it'd be y=2x + 1/4 ?.. i think All i need to know is the equation.

OpenStudy (anonymous):

y = 2x + b; now you will have to solve for b using the point (1,4) can you substitute x and y in the equation

OpenStudy (anonymous):

@JessNicole u there

OpenStudy (anonymous):

Yess, but your not understanding, It doesn't need solving. Slope=2, and the point (1,4) , these are the answers. So you'd graph (0,2) and (1,4) But they wanna know What is the equation for all this. :/

OpenStudy (anonymous):

thats what you will know when you solve for b you have the point (1,4) what is x and y in that point?

OpenStudy (anonymous):

I'm sorry, I really suck at math!!!:( So i should i substitue x for 1 and y for 4 ??

OpenStudy (anonymous):

y = 2*x + b correct now you have the point (1,4) 4 = 2*1 + b ----solve for b -2 -2 2 = b now you get the equation y = 2x + 2

OpenStudy (anonymous):

Ohh Okayy, THANK YOU!!!!! so much, i appreciate it:)